# Minimal Perimeter Triangle in Nonconvex Quadrilateral: Generalized Fagnano Problem

### v1

@@ 4 Solution of the Generalized Fagnano’s Problem:
Figure 14: Illustration of Case 6.1
Case 6.2
Suppose that $X$ lies on $AB$, $Y$ lies on $WD$, and $Z$ lies on $BC$.
Join $XW$. Since $\angle BWY>90^{\circ}$,in $\triangle XWY$ we have $XY>XW$, and in $\triangle ZWY$, we have $ZY>ZW$. Therefore,
$\text{per}(\triangle XZW)$ $\geq p_{W},$ (11)
$\text{per}(\triangle XZW)$ $=XZ+ZW+XW$
$\leq XZ+ZY+XY$
$\leq\text{per}(\triangle XYZ).$ (12)
Figure 15: Illustration of Case 6.2
Case 6.3
Suppose that $X$ lies on $AB$, $Y$ lies on $AW$, and $Z$ lies on $CD$. Let $XZ$ intersects $BW$ at $S$. Join $YS$. Since $\triangle XYS$ is inside the $\triangle XYZ$ with same base $XY$, Using Fact 2, we have $p_{W}\leq\text{per}(\triangle XYS)\leq\triangle XYZ$. Hence result follows.
Figure 16: Illustration of Case 6.3
Case 6.4
Suppose that $X$ lies on $AB$, $Y$ lies on $WD$, and $Z$ lies on $CD$.
Let $XZ$ intersects $BW$ at $S$. Join $SY$, In $\triangle SYW$ since $\angle SWY(=\angle BWY)>90^{\circ}$, therefore $SY>SW$. In $\triangle SZY,~SZ+ZY\geq SY$. Join $XW$. In $\triangle XYW$, $\angle XWY>90^{\circ}$ whereby $XY\geq WX$.
$p_{W}\leq\text{per}(\triangle XSW)$ $=XS+SW+WX$
$\leq XS+SY+WX$
$\leq XS+SZ+ZY+WX$
$\leq XZ+ZY+XY$
$\leq\text{per}(\triangle XYZ).$ (13)
Figure 17: Illustration of Case 6.4
Thus, in all Subcases 6.1, 6.2, 6.3 and 6.4, the orthic triangle of the auxiliary triangle $\triangle ABW$ is the unique minimal-perimeter triangle for Case (B) of Theorem 4.1.
The analyses of Cases (A) and (B) together establish that in every admissible configuration of an acute-angled nonconvex quadrilateral, the triangle of minimal perimeter is realized as the orthic triangle of an appropriately constructed auxiliary triangle—either $\triangle AMN$ or $\triangle ABW$. This unified result extends the classical Fagnano problem to nonconvex quadrilaterals and reveals a consistent geometric principle underlying both cases.
@@
It follows from Lemma 4.2 that the perimeter of the minimal triangle in the Generalized Fagnano Problem is bounded above by $2MN$ (or $2BW$ in Case-B), where $MN\perp AC$ in the nonconvex quadrilateral $ABCD$. This result establishes a simple and practical upper bound: the minimal perimeter cannot exceed twice the length of $MN$.
It may be noted that the authors did not find this generalization in the extensive literature on the subject, and it can be included as a natural and enriching exercise in textbooks that discuss the classical Fagnano problem.
@@ Acknowledgment
The second author gratefully acknowledges his supervisor, Dr. R. K. Pandey, for providing the opportunity to pursue this research and for his continuous mentorship and support throughout the course of this work.

### v2

@@ 4 Solution of the Generalized Fagnano Problem
Figure 11: Illustration of Case 5.1
Case 5.2
Suppose that $X$ lies on $AB$, $Y$ lies between $WD$, and $Z$ lies on $BC$.
Join $XW$. Since $\angle BWY>90^{\circ}$, in $\triangle XWY$ we have $|XY|>|XW|$, and in $\triangle ZWY$, we have $|ZY|>|ZW|$ (see figure 12). Therefore,
$p_{W}<\text{per}(\triangle XZW)$ $=|XZ|+|ZW|+|XW|$
$<|XZ|+|ZY|+|XY|$
$<\text{per}(\triangle XYZ).$ (9)
Figure 12: Illustration of Case 5.2
Case 5.3
Suppose that $X$ lies on $AB$, $Y$ lies on $AW$, and $Z$ lies between $CD$. Let $XZ$ intersects $BW$ at $S$. Join $YS$. Since $\triangle XYS$ is inside the $\triangle XYZ$ with same base $XY$, Using Fact 2, we have $p_{W}\leq\text{per}(\triangle XYS)<\text{per}(\triangle XYZ)$ (see figure 13). Hence result follows.
Figure 13: Illustration of Case 5.3
Case 5.4
Suppose that $X$ lies on $AB$, $Y$ lies between $WD$, and $Z$ lies on $CD$.
Let $XZ$ intersects $BW$ at $S$. Join $SY$, In $\triangle SYW$ since $\angle SWY(=\angle BWY)>90^{\circ}$, therefore $SY>SW$. In $\triangle SZY,~|SZ|+|ZY|>|SY|$. Join $XW$. In $\triangle XYW$, $\angle XWY>90^{\circ}$ whereby $|XY|>|WX|$ (see figure 14).
$p_{W}<\text{per}(\triangle XSW)$ $<|XS|+|SW|+|WX|$
$<|XS|+|SY|+|WX|$
$<|XS|+|SZ|+|ZY|+|WX|$
$<|XZ|+|ZY|+|XY|=\text{per}(\triangle XYZ)$
Figure 14: Illustration of Case 5.4
Thus, in all Subcases 5.1, 5.2, 5.3 and 5.4, the orthic triangle of the auxiliary triangle $\triangle ABW$ is the unique minimal-perimeter triangle for Case (B) of Theorem 4.1.
The analyses of Cases (A) and (B) together establish that in every admissible configuration of an acute-angled nonconvex quadrangle, the triangle of minimal perimeter is realized as the orthic triangle of an appropriately constructed auxiliary triangle—either $\triangle AMN$ or $\triangle ABW$. This unified result extends the classical Fagnano problem to nonconvex quadrangle and reveals a consistent geometric principle underlying both cases.
@@
It follows from Lemma 4.2 that the perimeter of the minimal triangle in the Generalized Fagnano Problem is bounded above by $2|MN|$ (or $2|BW|$ in Case-B), where $MN\perp AC$ in the nonconvex quadrangle $ABCD$. This result establishes a simple and practical upper bound: the minimal perimeter cannot exceed twice the length of $MN$ (or $2|BW|$).
@@ Acknowledgment
We sincerely thank the anonymous reviewer for the constructive comments and valuable suggestions, which greatly improved the quality and clarity of this manuscript. The second author gratefully acknowledges his supervisor, Dr. R. K. Pandey, for providing the opportunity to pursue this research and for his continuous mentorship and support throughout.
Disclosure Statement
The author declares that there is no conflict of interest regarding the publication of this paper.